An isolated solid sphere of aluminum having radius 7.0 cm is at a potential of 500 V. Calculate the number of electrons which have been removed from the sphere to raise it to this potential. 



1
Expert's answer
2021-06-02T09:34:31-0400

The potential and charge are


"V=\frac{kq}{r},\\\space\\
q=\frac{Vr}{k}."

The number of electrons:


"N=\frac qe=\frac{Vr}{ke}=2.43·10^{10}."